OpenStudy (anonymous):

solve for X : 8 over X equals X over 11?

OpenStudy (agreene):

x=+- 2*sqrt(22)

OpenStudy (anonymous):

x=sqrt(88)

OpenStudy (anonymous):

both of those answers are different options

OpenStudy (agreene):

8/x=x/11 8/x-x/11=0 (88-x^2)/11x=0 88-x^2=0 -x^2=-88 x^2=88 x=+-sqrt(88) x=+-2*sqrt(22) they are the same answer.

OpenStudy (anonymous):

if u want to confirm if they re right ,find the sqrt of 88 and substitute the value for x and check if the LHS of the equation equals the RHS